OpenStudy (anonymous):

Find three consecutive positive integers such that the sum of the squares of the first and second equals the square of the third.

OpenStudy (anonymous):

please help?

OpenStudy (lgbasallote):

let x = first integer x + 1 = second integer x + 2 = third integer sum of the squares of the first and second equals the square of the third so \[(x)^2 + (x+1)^2 = (x+2)^2\] \[\implies x^2 + x^2 + 2x + 1 = x^2 + 4x + 4\] \[ \implies 2x ^ 2 + 2x + 1 = x^2 + 4x + 4\] \[\implies 2x ^2 - x^2 + 2x - 4x + 1 - 4 = 0\] \[\implies x^2 - 2x - 3= 0\] \[\implies (x-3)(x+1) = 0\] therefore x = 3 and x = -1 does that help?

OpenStudy (anonymous):

yes thank you but, how did you get 2x from? and how are you combining x^2 and x^2 and getting 2x^2? can you explain please

OpenStudy (lgbasallote):

which 2x?

OpenStudy (anonymous):

in the second line of the equation

OpenStudy (anonymous):

i dont understand how you combined that and got a 2x

OpenStudy (lgbasallote):

\[x^2 + (x+1)^2\] (x+1)^2 = x^2 + 2x +1 therefore \[x^2 + (x+1)^2 = x^2 + x^2 + 2x + 1\] and then combine the two x^2 \[\implies 2x^2 + 2x +1\]

OpenStudy (anonymous):

If I might just point out that these are the sides of the classic 3:4:5 right triangle.

OpenStudy (lgbasallote):

maybe you're not familiar with the shortcut of (x+1)^2 i'll show it slowly \[(x+1)^2 \implies (x+1)(x+1) \implies x^2 + x + x + 1\] \[\implies x^2 + 2x +1\] that's where the 2x was coming from

OpenStudy (anonymous):

ohh thank you so so much! you're such a great help :)

OpenStudy (lgbasallote):

you're a great learner too ^_^